OpenStudy (anonymous):

A card is drawn from a standard 52 card deck. If the card is a king, you win $20; otherwise, you lose $3. What is the expected value of the game? a. Let x be the random variable for the amount won on a single play of this game. Complete the playoff table for this random variable X. Xi ___ ____ ( Maintain the order of values as given in the problem) Pi ___ ____ (type probabilities as simplified fractions)

OpenStudy (anonymous):

@satellite73 help mee

OpenStudy (anonymous):

\(X_1=20,X_2=-6\)

OpenStudy (anonymous):

\(P(X=20)=\frac{4}{52}=\frac{1}{13}\) and \[P(X=-6)=\frac{12}{13}\]

OpenStudy (anonymous):

\[E(X)=20\times \frac{1}{13}-6\times \frac{12}{13}=\frac{20-72}{13}\]

OpenStudy (anonymous):

so what is the answer of E(X)

OpenStudy (anonymous):

the one i wrote above, after you compute

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

-52/13

OpenStudy (anonymous):

so E(X)= -4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Or is the answer 14.46?

OpenStudy (anonymous):

no it is \(-4\)

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i have one lat question that i need help with quickly